OpenStudy (anonymous):

Three men walk inside a hotel and decide to stay for the night. They pay the price of $30 by splitting it so they each pay $10. They go up their room, but then the manager realizes that the price that night is $25 dollars so he sends $5 dollars up with the bell boy. The bellboy realizes that $5 can't be split by 3 so he pockets $2 and gives the men $1 each. But if you calculate it, the 3 men each payed $9, and 9 times 3 is 27, then add the $2 from the bellboy you got $29. Where did the last $1 go?

OpenStudy (anonymous):

Last 1 dollar is with the manager

OpenStudy (anonymous):

im sorry but its wrong

OpenStudy (anonymous):

Hmm good question.

OpenStudy (anonymous):

No idea

OpenStudy (anonymous):

30 - 5 = 25 + 2 = 27 / 3 = 9 25 was paid in... 3 was returned... 2 was pocketed 25 + 3 + 2 = 30 No money was missing.

OpenStudy (anonymous):

oHhh Nice

OpenStudy (nottim):

So the final information was really a red herring.

OpenStudy (anonymous):

lol

OpenStudy (phi):

You could say each paid $9, or $27. = $2 for the bellhop and $25 for the hotel. notice we do not add the $2 to the 27 (trying to get $30)l, we subtract $2 to get $25